OpenStudy (anonymous):

Identify the factors of x^2 - 5x - 24. (x + 8)(x - 3) (x - 8)(x + 3) (x + 4)(x - 6) (x - 4)(x + 6) @AravindG @marissalovescats @Whitemonsterbunny17

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

is it D

OpenStudy (anonymous):

The second one. Let me explain.

OpenStudy (anonymous):

wtf? where did you come from

OpenStudy (anonymous):

When you use the FOIL method the second one comes out to the expression

OpenStudy (anonymous):

its B dear...(x-8)(x+3)=x^2-8x+3x-24=x^2-5x-24

OpenStudy (anonymous):

oh thx ;) wow you guys are amazing

OpenStudy (anonymous):

@gregmb1999

OpenStudy (anonymous):

hey why'da block me